An outside loudspeaker (considered a small source) emits sound waves with a power output of 111 W (a) Find the intensity 10.7 m from the source. 0.76 Your response differs from the correct answer by more than 100%, w/m? (b) Find the intensity level in decibels at that distance. 118.8 Your response is within 10% of the correct value. This may be due to roundoff error, or you could have a mistake in your calculation Cary out all inter to at least four-digit accuracy to minimize roundoff error. dB edate results (c) At what distance would you experience the sound at the threshold of pain, 120 dB 2.97 Noed Help?4 11. 0-2 points SerCP10 14 P023 Your Teache A commuter train passes a passenger platform at a constant speed of 42.1 m/s. The train horn is sounded at its characteristic frequency of 330 Hz. (a) What overall change in frequency is detected by a person on the platform as the train moves from approaching to receding? (Indicate whether the frequency increases of decreases with the sign of your answer.) Hz (b) What wavelength is detected by a person on the platform as the train approaches?
Answer #1

a) we know the formula

intensity=P/4pir2=111W/4*pi*(10.7m)2=0.0772 W/m2

so the answer is 0.0772 W/m2 0.077 W/m2 or 0.08 W/m2

b) we have the formula

\beta=10 log(I/Io)=10 log( 7.72*10-2/1*10-12)=109 dB

so the answer is 109 dB or 108.88 dB or 108.9 dB

c) we have the formula

at 120 dB I=1 W/m2

I=p/4pir2

r=\sqrt{}P/4piI=\sqrt{}111/4pi*1=2.97m

answer 11) a) we know the formula

fo=fs(v/v-vs)

while approaching the speed=42.1 m/s=vs

fo=330m/s(343m/sm/s/343-42.1m/s)=376.17 Hz

when receding speed vs=-42.1m/s

fo=330m/s/(343m/s/343m/s-(-42.1m/s)=293.92Hz

so frequency shift is =f receding-f approach=293.92-376.17=-82 Hz

so the answer is -82 Hz or -82.25 Hz or -82.3 Hz.

b) we know the formula

\lambda=v/f( approach)=343m/s/376=0.912m

so the answer 0.912m or 0.91 m